# Non-Convex Over-the-Air Heterogeneous Federated Learning: A Bias–Variance Trade-off

### v3

@@ III Convergence Analysis and pre-scaler design
In this section, we theoretically characterize the learning performance of the proposed OTA-FL scheme as a function of the device pre-scalers $\{\gamma_{m}\}$. Since the global objective $F(\mathbf{w})$ is typically non-convex, we study convergence in terms of stationarity. To this end, our metric of interest is the expected (finite-time) average squared gradient norm of the global objective, given by $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\left\|\nabla F(\mathbf{w}_{t})\right\|^{2}]$. To study the convergence, we make the following standard assumptions:
@@
The sample-wise loss gradient for any given individual data sample $\xi$ is uniformly bounded, i.e., $\|\nabla\phi(\mathbf{w},\xi)\|\leq G_{\text{max}},\forall\mathbf{w}\in\mathbb{R}^{d}$. It follows from the triangle inequality that $\|\mathbf{g}_{m,t}\|\leq G_{\text{max}},\forall m,t$.
@@
The variance of local gradients to global gradient is bounded, that is, there exists $\kappa>0$ such that for all $\mathbf{w}\in\mathbb{R}^{d}\,,\,\frac{1}{N}\sum_{m\in[N]}\|\nabla f_{m}(\mathbf{w})-\nabla F(\mathbf{w})\|^{2}\leq\kappa^{2}$. Under Assumption 2, it further follows that $\kappa\leq 2G_{\text{max}}$.
Assumptions 1 (smoothness) and 3 are standard in FL analyses (e.g., [niid_fedavg, OTA_FL_H_data, OTA_FL_Optimization]), while Assumption 2 is also commonly adopted by many wireless FL works, see e.g., [Transmission_Power_Control_OTA_FA, OTA_FL_H_data, OTA_FL_Optimization]. The lower-boundedness in Assumption 1 is mild and typical for non-convex objectives [One_bit_FL, OTA_FL_Optimization]. Finally, Assumption 4, often called bounded gradient dissimilarity or data divergence, captures data heterogeneity and is widely used in the literature (e.g., [P_FL_bounded_niid]).
@@ III-A Main Convergence Results
Theorem 1 via (9) and (10) characterizes the convergence behavior of the proposed OTA-FL scheme through three key components: (1) an optimization term which depends on the initial FL model $\mathbf{w}_{0}$ that decays as $O(1/(\eta T))$; (2) a noisy global gradient estimation variance term; and (3) a model bias term induced by non-uniform average participation. Moreover, the variance term $\zeta$ splits further into: (a) transmission variance arising from truncated channel inversion, and intermittent devices transmission of the local gradients as a result of channel gain fluctuations, which randomizes device participation at each round; (b) mini-batch variance from stochastic mini-batch gradients; and (c) additive receiver noise. We highlight that the choice of pre-scalers $\{\gamma_{m}\}$ affects all three key components in the variance term as well as governs the resultant average model bias, critically revealing the fundamental bias-variance trade-off at the heart of this work. In particular, choosing smaller $\gamma_{m}$ lowers transmission variance and the bias term but also reduces $\alpha$, amplifying the receiver-noise term. Conversely, increasing $\gamma_{m}$ can enlarge $\alpha$ and suppress receiver noise, but shifts $\{p_{m}\}$ away from uniform, introducing a non-zero model bias. This trade-off calls for a careful optimization of the design parameters, which we develop in the next section.
@@ III-B OTA pre-scalers design
corresponding to the convergence bound in Theorem 1 excluding the initialization term, which is independent of the device pre-scalers $\{\gamma_{m}\}$. One can verify that due to the complicated dependence of bias and variance terms on the choice of pre-scalers $\{\gamma_{m}\}$, (P1) is a non-convex optimization. Leveraging variable dependencies, the optimization problem can also be expressed as a minimization over the (coupled) variables $\mathcal{X}\triangleq\{\{\gamma_{m}\},\{p_{m}\},\alpha\}$, subject to the following additional coupling constraints: (i) $\gamma_{m}\exp\!\{-\tfrac{\gamma_{m}^{2}G_{\max}^{2}}{d\Lambda_{m}E_{s}}\}=\alpha p_{m}$ (i.e., $\alpha_{m}=\alpha p_{m}$); (ii) $0<\gamma_{m}\leq\gamma_{m,\max}$, with $\gamma_{m,\max}=\sqrt{\tfrac{d\Lambda_{m}E_{s}}{2G_{\max}^{2}}}$; (iii) $0\leq\alpha\leq\min_{n}\frac{\alpha_{n,\max}}{p_{n}}$, with $\alpha_{m,\max}=\sqrt{\tfrac{d\Lambda_{m}E_{s}}{2e\,G_{\max}^{2}}}$; (iv) $p_{m}\in[0,1]$ and $\sum_{m}p_{m}=1$ (probability simplex). Note that $\alpha_{m}(\gamma_{m})$ is quasi-concave and peaks at $\gamma_{m,\max}$ with value $\alpha_{m,\max}$, thus (iii) enforces feasibility of (i) by ensuring $\alpha p_{m}=\alpha_{m}\leq\alpha_{m,\max}$ for every $m$. Moreover, for fixed $(\alpha,p_{m})$ the transmission-variance term increases with $\gamma_{m}$, and (i) admits two roots $\gamma_{m,1}\leq\gamma_{m,\max}\leq\gamma_{m,2}$, therefore the constraint (ii) causes no loss of optimality. Note that this optimization problem remains non-convex with non-convex terms present in both the objective and constraints. To address this, we adopt a successive convex approximation (SCA) approach, which iteratively solves convex surrogates of the original problem by linearizing the non-convex components around the current iterate. This process is guaranteed to converge to a stationary point of the original non-convex problem [Inner_approx].
@@
Each subproblem in (11) is convex and can be solved efficiently (e.g., via CVX [cvx]). Updating the anchors with the optimizer and iterating yields a high-quality solution to the original non-convex problem.
@@ IV Numerical Results
In this section, we evaluate the proposed SCA-optimized OTA-FL design. To this end, we study the popular handwritten digit classification problem in an FL setting on the widely used MNIST dataset [MNIST], which consists of C = 10 classes from “0” to “9” with each image of size 28 x 28 pixels. We performed experiments on a three-layer neural network architecture with ReLU non-linear activations and a hidden layer dimension of 1024, yielding a d=814,090d=814{,}090-dimensional parameter vector $\mathbf{w}$. The FL training uses an $\ell_{2}$-regularized cross-entropy loss at each device, with a regularization coefficient 0.01. We highlight that this setup results in a non-convex OTA-FL problem due to the choice of our model architecture. We consider the FL problem with $N=10$ devices uniformly deployed within a radius of $r_{\text{max}}=1750$ m from the PS situated at the center. The communication bandwidth is $B$ = 1 MHz with carrier frequency $f_{c}=2.4$ GHz, and the transmission power is set to $P_{\text{tx}}=$ 0 dBm. The noise power spectral density at the PS is $N_{0}=-173$ dBmW/Hz. The large-scale channel condition $\Lambda_{m}$ follows the log-distance path loss model, with a path loss exponent 2.2 and 50 dB loss at the reference distance of $1$ m. To simulate a challenging FL scenario, we adopt a heterogeneous (non-i.i.d.) data distribution. From a modified dataset of 10000 total samples (1000 per class), we adopt a data partition in which each device holds samples of exactly two digits (labels). Moreover, for any given label, its samples appear in the local datasets of no more than two devices. Each device uses full-batch gradients in every round, i.e., $|\mathcal{B}_{m,t}|=|\mathcal{D}_{m}|=1000$, which removes mini-batch variance in our simulations ($\sigma_{m}^{2}=0$ for all $m\in[N]$).
@@
We evaluate the performance of the proposed SCA-optimized OTA-FL design against five OTA-FL schemes from the existing literature. From [Opt_Power_Control_OTA_Comp], we include Optimized Power Control (OPC) OTA-Comp, which minimizes the per-round mean squared error (MSE) using global instantaneous CSI, and its low-complexity variant, LCPC OTA-Comp, which uses truncated channel inversion with a common pre-scaler optimized with statistical CSI. We also compare against the classical Vanilla OTA-FL [OTA_FL], which enforces zero instantaneous bias via channel inversion, requiring global instantaneous CSI. Finally, we include two heterogeneity-aware schedulers from [BB_FL]: BB-FL Interior, which schedules only devices within a fixed radius $R_{\text{in}}$, and BB-FL Alternative, which randomly alternates between full and interior scheduling. Obtaining global instantaneous CSI at the PS, as required by OPC OTA-Comp, Vanilla OTA-FL, and the BB-FL schemes, incurs substantial overhead. In contrast, our SCA design uses only statistical CSI at the PS together with local instantaneous CSI at the devices. We set $R_{\text{in}}=0.6\,r_{\max}$ for BB-FL methods (as in [BB_FL]), set $G_{\text{max}}=10$, and select the best constant stepsize $\eta$ for all schemes via grid search. For reference, we also report Ideal FedAvg, which performs the noiseless update in (2).
@@
Fig. 2 compares all methods over a fixed deployment, averaged across random fading and noise realizations. Fig. 2a reports the test accuracy versus rounds, while Fig. 2b shows the global objective. Ideal FedAvg demonstrates the best performance, since it aggregates gradients noiselessly. Among the practical wireless schemes, OPC OTA-Comp achieves the fastest convergence, closely followed by our proposed design, thanks to its optimized bias–variance trade-off. The strong performance of OPC OTA-Comp, however, comes at a substantial cost: it requires global instantaneous CSI at the PS in every round to optimize device pre-scalers. The proposed design, requiring only statistical CSI at the PS and local instantaneous CSI at devices, closely tracks OPC OTA-Comp in both metrics and attains competitive final accuracy, confirming that reducing model update variance while allowing a small average bias can yield performance gains.
@@
Compared with other baselines, the proposed design, despite lacking global CSI, consistently outperforms Vanilla OTA-FL, which enforces zero instantaneous bias at the cost of larger model update variance. While LCPC OTA-Comp also uses statistical CSI, it introduces an uncontrolled bias, leading to noticeably noisier and slower convergence. Finally, the wireless heterogeneity-aware schedulers of [BB_FL] behave as expected: BB-FL Alternative outperforms BB-FL Interior because restricting participation to a fixed interior subset reduces class coverage under our non-i.i.d. data split and weakens the model’s generalization. Overall, by optimizing the bias–variance trade-off, the proposed design achieves faster loss decay and stronger generalization performance than the baselines, without the global instantaneous CSI requirement.
@@ V Conclusion
In this paper, we have studied non-convex OTA-FL under heterogeneous wireless conditions. Departing from conventional zero-bias model updates, we proposed a novel OTA aggregation scheme that introduces a structured, time-invariant model bias to mitigate the high update variance caused by channel disparities. We established convergence to a stationary point and derived a bound, which explicitly exposes a bias-variance trade-off driven by device power control design. Guided by this theoretical insight, we formulated a non-convex joint power-control design and developed an efficient SCA algorithm that requires only statistical CSI at the base station. Numerical results on a non-convex image classification task validated our theory, demonstrating that under wireless heterogeneity, allowing a controlled bias can noticeably accelerate model convergence and improve model generalization relative to existing state-of-the-art OTA-FL baselines. Proof sketch of stationarity bound in (9): We start with $\nabla F(\mathbf{w}_{t}){=}\nabla\tilde{F}(\mathbf{w}_{t}){+}(\nabla F(\mathbf{w}_{t}){-}\nabla\tilde{F}(\mathbf{w}_{t}))$. Next, using $\|\mathbf{a}+\mathbf{b}\|^{2}\leq 2\|\mathbf{a}\|^{2}+2\|\mathbf{b}\|^{2}$ for arbitrary vectors $\mathbf{a},\mathbf{b}$, it follows that $\|\nabla F(\mathbf{w}_{t})\|^{2}\leq 2\|\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}+2\|\nabla F(\mathbf{w}_{t})-\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}.$ Hence, calculating the expectation and adding over $t$ yields $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{t})\|^{2}]$
@@
Recall that the model updates are given by $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta\hat{\boldsymbol{g}}_{t}$, where $\hat{\boldsymbol{g}}_{t}$ is the estimate of the global gradient, with expected value $\mathbb{E}[\hat{\boldsymbol{g}}_{t}|\mathbf{w}_{t}]=\nabla\tilde{F}(\mathbf{w}_{t})$, and variance bounded by $\mathbb{E}[\|\hat{\boldsymbol{g}}_{t}-\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]\leq\zeta$, given in (10). The detailed steps for obtaining this variance bound, which utilize Assumptions 2 and 3, are omitted here due to space constraints and can be found in our extended work [Biased_FL_Jorunal, Lemma 1]. It then follows that $\mathbb{E}[\mathbf{w}_{t+1}\!-\!\mathbf{w}_{t}|\mathbf{w}_{t}]=-\eta\nabla\tilde{F}(\mathbf{w}_{t})$ and $\mathbb{E}[\|\mathbf{w}_{t+1}\!-\!\mathbf{w}_{t}\|^{2}|\mathbf{w}_{t}]\leq\eta^{2}\|\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}+\eta^{2}\zeta$. Applying expectation conditional on $\mathbf{w}_{t}$ to both sides on (13), we further get

### v4

@@ III Convergence Analysis and pre-scaler design
In this section, we theoretically characterize the learning performance of the proposed OTA-FL scheme as governed by the choice of the device pre-scalers $\{\gamma_{m}\}$. In the non-convex setting considered here, we study convergence in terms of stationarity. To this end, our metric of interest is the expected (finite-time) average squared gradient norm of the global objective, given by $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\left\|\nabla F(\mathbf{w}_{t})\right\|^{2}]$. To study the convergence, we make the following standard assumptions:
@@
The sample-wise loss gradient for any given individual data sample $\xi$ is uniformly bounded, i.e., $\|\nabla\phi(\mathbf{w},\xi)\|\leq G_{\text{max}},\forall\;\mathbf{w}\in\mathbb{R}^{d}$. It then follows from the triangle inequality that $\|\mathbf{g}_{m,t}\|\leq G_{\text{max}},\forall m,t$.
@@
The variance of local gradients with respect to the global gradient is bounded, that is, there exists $\kappa>0$ such that for all $\mathbf{w}\in\mathbb{R}^{d}\,,\,\frac{1}{N}\sum_{m\in[N]}\|\nabla f_{m}(\mathbf{w})-\nabla F(\mathbf{w})\|^{2}\leq\kappa^{2}$. Under Assumption 2, it further follows that $\kappa\leq 2G_{\text{max}}$.
Assumptions 1 (smoothness) and 3 are standard in FL analyses (e.g., [9, 16, 25]), while Assumption 2 is also commonly adopted by many wireless FL works, see e.g., [4, 16, 25]. The lower-boundedness in Assumption 1 is mild and typical for non-convex objectives [23, 25]. Finally, Assumption 4, often called bounded gradient dissimilarity or data divergence, captures data heterogeneity and is widely used in the literature (e.g., [11]).
@@ III-A Main Convergence Results
Theorem 1 via (9) and (10) characterizes the convergence behavior of the proposed OTA-FL scheme through three key components: (1) an optimization term which depends on the initial FL model $\mathbf{w}_{0}$ that decays as $O(1/(\eta T))$; (2) a noisy global gradient estimation variance term; and (3) a model bias term induced by possibly non-uniform average participation. The variance term $\zeta$ splits further into: (a) transmission variance arising from truncated channel inversion, and intermittent gradient transmissions caused by channel fluctuations, which randomizes device participation at each round; (b) mini-batch variance from stochastic mini-batch gradients; and (c) additive receiver noise. Importantly, the choice of pre-scalers $\{\gamma_{m}\}$ affects all three key components in the variance term as well as governs the resultant average model bias, critically revealing the fundamental bias-variance trade-off at the heart of this work. In particular, a smaller $\gamma_{m}$ choice lowers transmission variance and the bias term but also reduces $\alpha$, amplifying the receiver-noise term. Conversely, increasing $\gamma_{m}$ can enlarge $\alpha$ and suppress receiver noise, but shifts $\{p_{m}\}$ away from uniform, introducing a non-zero model bias. This trade-off calls for a careful optimization of the design parameters, which we develop in the next section.
@@ III-B OTA pre-scalers design
corresponding to the convergence bound in Theorem 1 excluding the initialization term, which is independent of the device pre-scalers $\{\gamma_{m}\}$. One can verify that due to the complicated dependence of bias and variance terms on the choice of pre-scalers $\{\gamma_{m}\}$, (P1) is a non-convex optimization. Leveraging variable dependencies, we equivalently rewrite the problem as a minimization over the coupled variables $\mathcal{X}\triangleq\{\{\gamma_{m}\},\{p_{m}\},\alpha\}$, subject to the following additional coupling constraints: (i) $\gamma_{m}\exp\{-\tfrac{\gamma_{m}^{2}G_{\max}^{2}}{d\Lambda_{m}E_{s}}\}=\alpha p_{m}$ (i.e., $\alpha_{m}=\alpha p_{m}$); (ii) $0<\gamma_{m}\leq\gamma_{m,\max}$, with $\gamma_{m,\max}=\sqrt{\tfrac{d\Lambda_{m}E_{s}}{2G_{\max}^{2}}}$; (iii) $0\leq\alpha\leq\min_{n}\frac{\alpha_{n,\max}}{p_{n}}$, with $\alpha_{m,\max}=\sqrt{\tfrac{d\Lambda_{m}E_{s}}{2e\,G_{\max}^{2}}}$; (iv) $p_{m}\in[0,1]$ and $\sum_{m\in[N]}p_{m}=1$ (probability simplex). Since $\alpha_{m}(\gamma_{m})$ is quasi-concave and attains its maximum $\alpha_{m,\max}$ at $\gamma_{m,\max}$, the constraint (iii) enforces feasibility of (i) by ensuring $\alpha p_{m}=\alpha_{m}\leq\alpha_{m,\max}$ for all $m$. Moreover, for fixed $(\alpha,p_{m})$ the transmission-variance term increases with $\gamma_{m}$, and (i) admits two roots $\gamma_{m,1}\leq\gamma_{m,\max}\leq\gamma_{m,2}$, therefore the constraint (ii) causes no loss of optimality. Notably, the resulting problem remains non-convex due to non-convex terms in both the objective and constraints. To address this, we adopt a successive convex approximation (SCA) approach, which iteratively solves convex surrogates of the original problem by linearizing the non-convex components around the current iterate. This procedure is guaranteed to converge to a stationary point of the original non-convex problem [12].
@@
Each subproblem in (11) is convex and can be solved efficiently (e.g., via CVX [7]). We initialize the anchors with a feasible point and iterate until the objective improvement falls below a tolerance threshold or a maximum number of iterations is reached, yielding a high-quality solution to the original non-convex problem.
@@ IV Numerical Results
In this section, we evaluate the proposed SCA-optimized OTA-FL design. To this end, we study the popular handwritten digit classification problem in an FL setting on the widely used MNIST dataset [8], which consists of C = 10 classes from “0” to “9” with each image of size 28 x 28 pixels. We performed experiments on a fully connected neural network with one hidden layer, ReLU non-linear activations, and a hidden layer dimension of 1024, yielding a d=814,090d=814{,}090-dimensional parameter vector $\mathbf{w}$. The FL training uses an $\ell_{2}$-regularized cross-entropy loss at each device, with a regularization coefficient of 0.01. The setup results in a non-convex OTA-FL problem. We consider the FL problem with $N=10$ devices uniformly deployed within a radius of $r_{\text{max}}=1750$ m from the PS situated at the center. The communication bandwidth is $B$ = 1 MHz with carrier frequency $f_{c}=2.4$ GHz, and the transmission power is set to $P_{\text{tx}}=$ 0 dBm. The noise power spectral density at the PS is $N_{0}=-173$ dBm/Hz. The large-scale channel condition $\Lambda_{m}$ follows the log-distance path loss model, with a path loss exponent 2.2 and 50 dB loss at the reference distance of $1$ m. To simulate a challenging FL scenario, we consider a heterogeneous (non-i.i.d.) data distribution. From a modified dataset of 10000 total samples (1000 per class), we adopt a data partition in which each device holds samples of exactly two digits (labels). Moreover, for any given label, its samples appear in the local datasets of no more than two devices. Each device computes full-batch gradients in every round, i.e., $|\mathcal{B}_{m,t}|=|\mathcal{D}_{m}|=1000$, which removes mini-batch variance in our simulations ($\sigma_{m}^{2}=0$ for all $m\in[N]$).
@@
We evaluate the performance of the proposed SCA-optimized OTA-FL design against five OTA-FL baselines. From [5], we include Optimized Power Control (OPC) OTA-Comp, which minimizes the per-round mean squared error (MSE) using global instantaneous CSI, and its low-complexity variant, LCPC OTA-Comp, which uses truncated channel inversion with a common pre-scaler optimized with statistical CSI. We also compare against the classical Vanilla OTA-FL [21], which enforces zero instantaneous bias via channel inversion, requiring global instantaneous CSI. Finally, we include two heterogeneity-aware schedulers from [24]: BB-FL Interior, which schedules only devices within a fixed radius $R_{\text{in}}$, and BB-FL Alternative, which randomly alternates between full and interior scheduling. Acquiring global instantaneous CSI at the PS, as required by OPC OTA-Comp, Vanilla OTA-FL, and the BB-FL schemes, incurs substantial overhead. In contrast, our SCA design uses only statistical CSI at the PS together with local instantaneous CSI at the devices. We set $R_{\text{in}}=0.6\,r_{\max}$ for BB-FL methods (as in [24]), set $G_{\text{max}}=10$, and select the best constant stepsize $\eta$ for all schemes via grid search. For reference, we also report Ideal FedAvg, which performs the noiseless update in (2).
@@
Fig. 2 compares all methods over a fixed deployment, averaged across random fading and noise realizations. Fig. 2a reports the test accuracy versus rounds, while Fig. 2b shows the global objective. Ideal FedAvg demonstrates the best performance, since it aggregates gradients noiselessly. Among the practical wireless schemes, OPC OTA-Comp achieves the fastest convergence, closely followed by our proposed design, thanks to its optimized bias-variance trade-off. The strong performance of OPC OTA-Comp, however, comes at a substantial cost: it requires global instantaneous CSI at the PS in every round to optimize device pre-scalers. The proposed design, requiring only statistical CSI at the PS and local instantaneous CSI at devices, closely tracks OPC OTA-Comp in both metrics and attains competitive final accuracy, confirming that reducing model update variance while allowing a small average bias can yield performance gains.
@@
Compared with other baselines, the proposed design, despite lacking global CSI, consistently outperforms Vanilla OTA-FL, which enforces zero instantaneous bias at the cost of larger model update variance. While LCPC OTA-Comp also uses statistical CSI, it can introduce a less controllable bias, leading to noticeably noisier and slower convergence. Finally, the wireless heterogeneity-aware schedulers of [24] behave as expected: BB-FL Alternative outperforms BB-FL Interior because restricting participation to a fixed interior subset reduces class coverage under our non-i.i.d. data split and weakens the model’s generalization. Overall, by optimizing the bias-variance trade-off, the proposed design achieves faster loss decay and stronger generalization than the baselines without requiring global instantaneous CSI.
@@ V Conclusion
In this paper, we have studied non-convex OTA-FL under heterogeneous wireless conditions. Departing from conventional zero-bias model updates, we proposed an OTA aggregation scheme that introduces a structured, time-invariant model bias to mitigate the high update variance caused by channel disparities. We established convergence to a stationary point and derived a bound, which explicitly exposes a bias-variance trade-off driven by device power control design. Guided by this theoretical insight, we formulated a non-convex joint power-control design and developed an efficient SCA algorithm that requires only statistical CSI at the base station. Numerical results on a non-convex image classification task validated our theory, demonstrating that under wireless heterogeneity, allowing a controlled bias can noticeably accelerate model convergence and improve model generalization relative to state-of-the-art OTA-FL baselines. Proof sketch for (9):
We start with $\nabla F(\mathbf{w}_{t}){=}\nabla\tilde{F}(\mathbf{w}_{t}){+}(\nabla F(\mathbf{w}_{t}){-}\nabla\tilde{F}(\mathbf{w}_{t}))$. Next, using $\|\mathbf{a}+\mathbf{b}\|^{2}\leq 2\|\mathbf{a}\|^{2}+2\|\mathbf{b}\|^{2}$ for arbitrary vectors $\mathbf{a},\mathbf{b}$, it follows that $\|\nabla F(\mathbf{w}_{t})\|^{2}\leq 2\|\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}+2\|\nabla F(\mathbf{w}_{t})-\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}.$ Hence, taking expectations and averaging over $t$ yields $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{t})\|^{2}]$
@@
Recall that the model updates are given by $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta\hat{\boldsymbol{g}}_{t}$, where $\hat{\boldsymbol{g}}_{t}$ is the estimate of the global gradient, with expected value $\mathbb{E}[\hat{\boldsymbol{g}}_{t}|\mathbf{w}_{t}]=\nabla\tilde{F}(\mathbf{w}_{t})$, and variance bounded by $\mathbb{E}[\|\hat{\boldsymbol{g}}_{t}-\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]\leq\zeta$, where $\zeta$ is given in (10). The derivation of this variance bound (using Assumptions 23) is omitted for brevity and can be found in our extended work [2, Lemma 1]. It then follows that $\mathbb{E}[\mathbf{w}_{t+1}\!-\!\mathbf{w}_{t}|\mathbf{w}_{t}]=-\eta\nabla\tilde{F}(\mathbf{w}_{t})$ and $\mathbb{E}[\|\mathbf{w}_{t+1}\!-\!\mathbf{w}_{t}\|^{2}|\mathbf{w}_{t}]\leq\eta^{2}\|\nabla\tilde{F}(\mathbf{w}_{t})\|^{2}+\eta^{2}\zeta$. Applying expectation conditional on $\mathbf{w}_{t}$ to both sides on (13), we get